The diameter of a pipe at the section 1 and 2 are 15 cm and 20 cm respectively. Find the discharge through the pipe if velocity of water at section 1 is 4 m/s. Determine also the velocity at section 2.
1
Expert's answer
2020-12-08T06:39:13-0500

Diameter of pipe 1 "d_1=15 cm=0.15\ m"

C/S Area of pipe 1 "A_1 =\dfrac{\pi(d_1)^2}{4}=0.005625\pi\ m^2"


Diameter of pipe 2 "d_2=20\ cm=0.2\ m"

C/S Area of pipe 2 "A_2 =\dfrac{\pi(d_1)^2}{4}=0.01\pi\ m^2"


Velocity at pipe 1 "v_1=4\ m/s"

Velocity at pipe 2 "v_2=?"


Discharge through pipe 1 "Q_1=A_1v_1=0.005625\pi\ m^2 (4\ m/s)"

"=0.0225\pi\ m^3/s=0.0707\ m^3/s"

"Q_1=0.0707\ m^3/s"


From continuity "Q_1=Q_2+Q_3"


Discharge through pipe 2 "Q_2=A_2v_2=0.01\pi\ m^2 (v_2)"

"=Q_1=0.0225\pi\ m^3/s"


Velocity at pipe 2 "v_2=\dfrac{Q_2}{A_2}=\dfrac{0.0225\pi\ m^3/s}{0.01\pi\ m^2 }=2.25\ m/s"

"v_2=2.25\ m/s"